# Star formation rate density as a function of galaxy mass at z 0.2 with MUSE and GAMA surveys

### v1

@@ 6 Discussion
The SFRD results from the combined sample using the MUSE Wide and MUSE HUDF catalogues, and the ones obtained using the catalogues from the GAMA survey, show the behaviour of the SFRD down to stellar masses of 105.5 M⊙ with MUSE and 106 M⊙ with GAMA, pushing our understanding of the SFRD to the faintest galaxies. The results from this work indicate that very low-mass galaxies have a significant contribution to the cosmic SFRD, although at lower levels than more massive galaxies. While individual low-mass galaxies have low SFRDs, their potentially high number density means they could contribute significantly to the total cosmic SFRD.
@@
The MUSE sample benefited from accurate H emission line flux measurements, thanks to MUSE Wide and MUSE HUDF advanced IFU spectroscopy, therefore the SFR for this sample was derived using the calibration described in Sec. 3.1. However, since the sample only contained a limited number of 27 galaxies, the SFRD could only be derived using large stellar mass bins of , in order to ensure a reasonable pseudo-Poisson error. Nonetheless, the results show reasonable agreement with the extrapolation based on previous studies, and with the SFRD estimated using GAMA.
@@
The SFRs used for the analysis of the GAMA sample were derived via SED fitting using MagPhys (Da Cunha et al., 2008), as the H emission line measurements for low-mass galaxies in GAMA might be inflated and too susceptible to corrections required for aperture effects. However, its large area allowed us to create a sample of 7579 star forming galaxies at z 0.06. As shown in Fig. 9, the slope of the GAMA SFRD is similar to the one found using MUSE at low masses. There is an increase in the SFRD derived from GAMA as stellar mass increases up to 1010.5 M⊙. There is a sharp decline at the higher end of the function as expected from the exponential cutoff in the GSMF and from the higher fraction of quenched galaxies at higher masses.
@@
Previous studies encountered difficulties in estimating the SFRD below M⊙, as dwarf galaxies are challenging to observe. Gilbank et al. (2010b) derived the SFRD down to stellar masses of M⊙ using H, [OII], and -band luminosities from the Sloan Digital Sky Survey (SDSS), whereas Sedgwick et al. (2019) managed to estimate the SFRD at 107 M⊙ using CCSNe as "signposts" to low surface brightness galaxies to constrain their abundance and contribution to the total star formation rate density (SFRD) at lower masses. As shown in Fig. 9, both studies suggested a constant decline in SRFD, similarly to the results found in this work.
A constant decline in SFRD at lower stellar masses could have great significance in our understanding of the GSMF and galaxy evolution. In fact, the results suggest that there is no turn-over at the faint end of the GSMF, suggesting that low-mass galaxies might be more abundant than previously thought. The implications of such findings would affect our understanding of galaxy evolution and highlight the importance of low-mass galaxies in various cosmological processes. In fact, the absence of a clear turn-over in the GSMF might indicate a higher number of low-mass dark matter halos than expected, posing serious questions to the current simulations of dark matter structure formation (Papastergis et al., 2011). Moreover, if the number of low-mass galaxies that are actively forming stars is higher than previously anticipated, it could also mean that faint galaxies played an important role during the epoch of reionization of the universe (Livermore et al., 2017).
@@
The volume of the MUSE sample is only 804 Mpc3. Therefore this is potentially subject to significant cosmic variance. To test this, we consider previous observations of the GSMF from the GAMA survey (Baldry et al., 2012; Wright et al., 2017). Measurement of the number density for galaxies more massive than 108 M⊙ are not affected by low-surface-brightness incompleteness in GAMA. Integrating the number density from the high-mass end, we obtain between 0.028 and 0.036 galaxies per Mpc3 by changing the log-mass limit between 7.9 to 8.1. This allows for some uncertainty in how stellar mass is estimated.
@@
As the volume of the combined MUSE sample is 804 Mpc3, the expected number of galaxies with stellar masses higher than 108 M⊙ should be between 22 and 29. However, the total number of galaxies above this mass limit, including passive non star-forming galaxies detected using AUTOZ (Baldry et al., 2014), in the volume of interest is only 12. These results suggest that the volume is underdense. If we assume that the number density of galaxies below 108 M⊙ scales in the same way, we can account for the underdensity by multiplying the values of the SFRD by about two.
@@
This simple correction can be seen in the plot reported in Fig. 10, where it can be seen how by accounting for underdensity, the SFRD is more consistent with the values derived in previous studies (Sedgwick et al., 2019; Gilbank et al., 2010b). These results suggest that a larger sample is required to have a conclusive understanding of the low-mass galaxy population and their impact onto the SFRD and GSMF. These results may also provide a useful benchmark for comparisons with future deep field surveys and cosmological simulations.
@@ 7 Summary
In summary, the results from this work show that:
The MUSE Wide and MUSE HUDF surveys have a total volume of 804 Mpc3. Their exceptional depth allowed us to detect 27 star-forming galaxies with H emission lines at a chosen inspected redshift of z 0.2. On the other hand, the sample from the GAMA survey benefited from a much larger number of galaxies, with a total of 7579 galaxies at redshift z 0.06. However, the H emission line flux measurements of the sample may be overestimated, due to the inaccuracies of fiber spectroscopy, and we had to rely on SED fitting to estimate accurate SFRs.
The SFRD with respect to stellar mass distribution was estimated for the MUSE Wide and HUDF combined sample, using SFR H measurements, and the GAMA sample, using SFRs values derived with MagPhys SED fitting. The results shown in Fig. 9 suggest that there is a constant slope in SFRD at lower masses with values of the slope (Eq. 12) between 0.57 to 0.73. This could indicate that very low-mass galaxies have a significant contribution to the cosmic SFRD.
The slope in the SFRD(M) implies that the faint-end slope in the GSMF, , is in the range to (see Eqs. 13–15 for this argument), consistent with values found in GAMA (Baldry et al., 2012; Wright et al., 2017), providing important evidence of the absence of a turn-over at the faint end of the GSMF, suggesting that there might be more low-mass galaxies than previously thought. This might be of huge significance for our understanding of galaxy formation and evolution.
When considering the number density of galaxies more massive than M⊙ using the GAMA GSMFs compared with the number of galaxies detected at in MUSE Wide and MUSE HUDF, it implies that the volume considered in MUSE is an underdense region, with about a factor of two fewer galaxies than expected.
The results could be improved with larger surveys, allowing us to investigate the faint end of the function, and with comparisons with cosmological simulations. For example, the Rubin observatory can be used to detect and characterize the SFRD using CCSNe; while Euclid plus redshifts from 4MOST provide a deeper galaxy redshift survey. And we can also use more MUSE fields (blind to low-redshift volume) for H selected galaxy samples.
@@ Acknowledgements
We would like to express my sincere gratitude to Prof. Phil James for his invaluable guidance and support throughout this work. We also thank the MUSE collaboration for providing the data used in this research, particularly from the MUSE WIDE and MUSE HUDF surveys. Finally, GM acknowledges the financial support from the Science and Technology Facilities Council (STFC) for funding his PhD.

### v2

@@ 6 Discussion
The SFRD results from the combined sample using the MUSE Wide and MUSE HUDF catalogues, and the results obtained using the catalogues from the GAMA survey, show the behaviour of the SFRD down to stellar masses of with MUSE and 106 Mā with GAMA, pushing our understanding of the SFRD to the faintest galaxies. The results from this work indicate that low-mass galaxies () contribute about 1% of the cosmic SFRD (per dex) relative to massive galaxies at the peak of the SFRD().
@@
The MUSE sample benefited from accurate H emission line flux measurements, thanks to MUSE Wide and MUSE HUDF advanced IFU spectroscopy, therefore the SFR for this sample was derived using the calibration described in Sec. 3.1. However, since the sample only contained a limited number of 27 galaxies, the SFRD could only be derived using large stellar mass bins of , in order to ensure a reasonable Poisson error. Nonetheless, the results show reasonable agreement with the extrapolation based on previous studies, and with the SFRD estimated using GAMA.
@@
The SFRs used for the analysis of the GAMA sample were derived via SED fitting using MagPhys (DaĀ Cunha etĀ al., 2008), as the H emission line measurements for low-mass galaxies in GAMA might be inflated and too susceptible to corrections required for aperture effects. However, its large area allowed us to create a sample of 7579 star forming galaxies at . As shown in Fig. 7, the slope of the GAMA SFRD is similar to the one found using MUSE at low masses. There is an increase in the SFRD derived from GAMA as stellar mass increases up to . There is a sharp decline at the higher end of the function as expected from the exponential cutoff in the GSMF and from the higher fraction of quenched galaxies at higher masses.
@@
Previous studies encountered difficulties in estimating the SFRD below , as dwarf galaxies are challenging to observe. Gilbank etĀ al. (2010b) derived the SFRD down to stellar masses of using H, [OII], and -band luminosities from the Sloan Digital Sky Survey (SDSS), whereas Sedgwick etĀ al. (2019) managed to estimate the SFRD at using CCSNe as "signposts" to low surface brightness galaxies to constrain their abundance and contribution to the total star formation rate density (SFRD) at lower masses. As shown in Fig. 7, both studies suggested a constant decline in SFRD, similarly to the results found in this work.
The SFRD measurements of Sedgwick etĀ al. (2019) and Gilbank etĀ al. (2010b) do appear offset from our work. However, Sedgwick etĀ al. used a factor of two correction at low galaxy masses for the visibility of CCSNe, and there is uncertainty over classification of CCSNe for some of their sample. This also requires a CCSNe rate to SFRD calibration. So this SFRD from CCSNe may be overestimated. Gilbank etĀ al. used a different calibration of H luminosity to SFR (from Kennicutt 1998). This SFRD is shown adjusted to our calibration () in Fig.Ā 8.
A constant decline in SFRD at lower stellar masses could have great significance in our understanding of the GSMF and galaxy evolution. In fact, the results () suggest that there is no turn-over at the faint end of the GSMF down to . This is based on the assumption that the main sequence slope (Eq.Ā 14) is approximately unity e.g.Ā Lee etĀ al. (2015). Or in other words, the mean specific SFR of star-forming galaxies is similar for mass ranges below (James etĀ al., 2008b). This is expected if these galaxies have formed their stars quasi-continuously since galaxy formation began.
A reasonable estimate of the SFRD does not require that we find all star-forming galaxies in a given volume. It is plausible that low-mass star-forming galaxies may go through lull periods (but not quenched) which would reduce their H luminosity significantly and thus not appear in H-selected samples. This is the duty cycle of star formation and it is likely more extreme at low masses. In effect, the SFRD over a cosmological volume is an estimate of the time average over a population of star-forming galaxies. This is because we do not expect the chance of being in a lull or burst phase being correlated with other star-forming dwarf galaxies.
The implications of our findings highlights the importance of low-mass galaxies in various cosmological processes. The slope of the SFRD suggests a steep faint-end slope for the GSMF, consistent with prior studies (e.g., Wright etĀ al. 2017). This highlights the importance of low-mass galaxies in the star formation budget and their potential role in early cosmic epochs.
@@
The volume of the MUSE sample is only 804 Mpc3. Therefore this is potentially subject to significant cosmic variance. To test this, we consider previous observations of the GSMF from the GAMA survey (Baldry etĀ al., 2012; Wright etĀ al., 2017). Measurements of the number density for galaxies more massive than are not affected by low-surface-brightness incompleteness in GAMA. Integrating the number density from the high-mass end, we obtain between 0.028 and 0.036 galaxies per Mpc3 by changing the log-mass limit between 7.9 to 8.1. This allows for some uncertainty in how stellar mass is estimated.
@@
As the volume of the combined MUSE sample is 804 Mpc3, the expected number of galaxies with stellar masses higher than should be between 22 and 29. However, the total number of galaxies above this mass limit, including passive (non star-forming) galaxies detected using AUTOZ (Baldry etĀ al., 2014), in the volume of interest is only 12. These results suggest that the volume is underdense. If we assume that the number density of galaxies below scales in the same way, we can account for the underdensity by multiplying the values of the SFRD by about two.
@@
This simple correction can be seen in the plot reported in Fig. 8, where it can be seen how by accounting for underdensity, the SFRD is more consistent with the values derived in previous studies (Sedgwick etĀ al., 2019; Gilbank etĀ al., 2010b). These results suggest that a larger sample is required to have a conclusive understanding of the low-mass galaxy population and their impact on the SFRD and GSMF. These results may also provide a useful benchmark for comparisons with future deep field surveys and cosmological simulations.
@@ 7 Summary
This study investigates the contribution of star-forming galaxies to the SFRD as a function of stellar mass at , using MUSE Wide, MUSE HUDF, and GAMA surveys. By combining deep spectroscopic coverage and wide-area data, we probe the faint end of the SFRD down to stellar masses of , providing new insights into the low-mass galaxy population. In summary, the results from this work show that:
ā¢
The MUSE Wide and MUSE HUDF surveys have a total volume of 804 Mpc3. Their exceptional depth allowed us to detect 27 star-forming galaxies with H emission lines at redshifts . On the other hand, the sample from the GAMA survey benefited from a much larger number of galaxies, with a total of 7579 galaxies at redshift . For GAMA, we rely on the photometric estimates of the SFRs from SED fitting because the spectroscopic H measurements require significant aperture corrections.
- ā¢
ā¢
The SFRD with respect to stellar mass distribution was estimated for the MUSE Wide and HUDF combined sample, using SFR H measurements, and the GAMA sample, using SFRs values derived with MagPhys SED fitting. The results shown in Fig. 7 suggest that there is a constant slope in SFRD at lower masses with values of the slope (Eq.Ā 12) between 0.57 to 0.73. This quantifies the contribution of low-mass galaxies to the cosmic SFRD.
- ā¢
ā¢
When considering the number density of galaxies more massive than using the GAMA GSMFs compared with the number of galaxies detected at in MUSE Wide and MUSE HUDF, there is a significant discrepancy between the number densities. This implies that the volume considered in MUSE is underdense, with about a factor-of-two fewer galaxies than the cosmic mean.
ā¢
The results could be improved with larger surveys, allowing us to more accurately investigate the faint end of the SFRD and GSMF, and with comparisons with cosmological simulations. For example, the Rubin observatory can be used to detect and characterize the SFRD using CCSNe; while Euclid imaging (to detect low surface-brightness galaxies) plus redshifts from 4MOST provide a deeper galaxy redshift survey. And we could also use more MUSE fields (blind to low-redshift volume) for H selected galaxy samples.
@@ Acknowledgements
We thank the anonymous referee for thoughtful comments that significantly improved the paper. We would like to express our sincere gratitude to Prof.Ā Phil James for his invaluable guidance and support throughout this work. We also thank the MUSE collaboration for providing the data used in this research, particularly from the MUSE WIDE and MUSE HUDF surveys. Finally, GM acknowledges the financial support from the Science and Technology Facilities Council (STFC) for funding his PhD.